OpenStudy (anonymous):

I can't have any guess about the answer to the question, I really need help so please no direct answers help me understand how to solve this. Will medal Muriel opened a savings account with an annual interest rate of 9% and an initial deposit of $4000. If her interest is compounded quarterly, how much is in her account after 6 years? Round your answer to the nearest cent. interest compounded quarterly: A = P (1 +r/4)4t Answer:_______

OpenStudy (welshfella):

the formula for compound interest paid quarterly is A = P ( 1 + r/4)^4n where A is the amount after n years , r = annual rate ( as a decimal) and n = number of years Oh - i 've just see you have that already!!!

OpenStudy (welshfella):

all you need do is plug in n = 6 , r = 0.09 ( which is same as 9%) , P = 4000

OpenStudy (anonymous):

@welshfella so is the answer 4000

OpenStudy (anonymous):

@Hayhayz @HaileyD @imqwerty @Preetha @Redcan

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

Ok A=4000(1+0.09/4)4*6 4000= principal Percent rate= 9%=0.09 t=time=6 Pretty sure that's how you set it up

OpenStudy (anonymous):

so the answer is 98160

OpenStudy (raffle_snaffle):

@welshfella is he right?

OpenStudy (welshfella):

i'll check now

OpenStudy (welshfella):

i get 6823.07

OpenStudy (anonymous):

I got 42,624.00

OpenStudy (anonymous):

I'll go with 6823.07, I really hope this is right

OpenStudy (raffle_snaffle):

OP another suggestion to solve is use i_a = r/m then you can use F = P*(1+i)^n but your n is in quarters instead of years.

OpenStudy (welshfella):

= 4000(1 + 0.0225)^24

OpenStudy (raffle_snaffle):

wait... We are solving for A or F? llol

OpenStudy (welshfella):

yea 6923.07

OpenStudy (anonymous):

Nerdykid, did welsh get it right?

OpenStudy (welshfella):

solving for the amount after 6 years

OpenStudy (raffle_snaffle):

Haha I meant A = P(A/P, i, 6)

OpenStudy (anonymous):

I don't know yet I'll tell you once I'm done

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Welsh can you help me with my question?